the price of a coffe and pastry is $7.75. the coffee cost $0.75 more than the pastry. what are the individual cost of one coffe and one pastry?
Answer by macston(5194)

P = cost of pastry; C = cost of coffee=P+$0.75
P+C=$7.75 Substitute for C
P+P+$0.75=$7.75 Subtract $0.75 from each side.
2P=$7.00 Divide each side by 2.
P=$3.50 ANSWER: A pastry costs $3.50
C=P+$0.75=$3.50+$0.75=$4.25 ANSWER: Coffee costs $4.25
CHECK:
P+C=$7.75
$3.50+$4.25=$7.75
$7.75=$7.75